Given the constraints
A+B + C ≤ 24, B +C ≥ 8 and A ≥ 0, B ≥ 0, C ≥ 0.
Maximize 24-A-B - C
A: amount of time spent on school work
B: amount of time spent on fun
C: amount of time spent on pay work
1
Expert's answer
2020-04-22T14:16:20-0400

"max\quad 24-A-B-C"

subject to,

"A+B + C \le24\\
B +C\ge8\\
A ≥ 0, B ≥ 0, C ≥ 0"


"max\quad 24-A-B-C+0S_1+0S_2-Ma_1"

subject to,

"A+B + C+S_1\hspace{2.1 em} =24\\
\hspace{2 em}B +C-S_2+a_1=8\\
A , B, C,S_1,S_2,a_1 ≥ 0"


Iteration-1


"a_1" is goin out from the basis and B is come into the basis.


Iteration-2


since all the "Z-C_j\ge0" optimum solution is obtain.


optimum solution is,

"\bold{A=0\quad B=8\quad C=0}\\
Z_{max}=24-8\\
\bold{Z_{max}=16}"